OpenStudy (anonymous):

PLEASE HELP!!!!!! A function is shown below: f(x) = x3 + 5x2 - x - 5 Part A: What are the factors of f(x)? Show your work.

OpenStudy (anonymous):

since \(f(1)=0\) this will factor as \((x-1)(something)\)

OpenStudy (anonymous):

ok, and?

OpenStudy (anonymous):

@satellite73 ?

OpenStudy (anonymous):

you have a couple of choices you can do long division (a pain ) you can use synthetic division (easiest) or you can think (requires thinking)

OpenStudy (anonymous):

i cannot write the first two methods here it is annoying we can use the think method if you like

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[x^3 + 5x^2 - x - 5=(x-1)(ax^2+bx+c)\] and \(a\) and \(c\) are pretty obvious right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what are a and c?

OpenStudy (anonymous):

ok lets go slow first we know that \(x-1\) is a factor because \(f(1)=0\) right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

that means \[x^3+5x^2-x-5=(x-1)(something)\]

OpenStudy (anonymous):

and the "something" has to be a quadratic (degree two) because when you multiply out you have to get a polynomial of degree three, namely \(x^3+5x^2-x-5\)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so what can that quadratic be?

OpenStudy (anonymous):

um, x^2 + 5x - 5?

OpenStudy (anonymous):

well i don't know maybe

OpenStudy (anonymous):

\[x^3 + 5x^2 - x - 5=(x-1)(ax^2+bx+c)\] is what i wrote and i said that \(a\) is obvious because it is \(a=1\) for sure because when you multiply out you have to get \(x^3\)

OpenStudy (anonymous):

and \(c=5\) not \(-5\) because when you multiply out you need to end up with \(-5\)

OpenStudy (anonymous):

right. so the factors are (x-1) (x + 1) and (x + 5) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

yw